# High-dimensional maximum-entropy phase space tomography using normalizing flows

### v1

@@ Abstract
Particle accelerators generate charged particle beams with tailored distributions in six-dimensional (6D) position-momentum space (phase space). Knowledge of the phase space distribution enables model-based beam optimization and control. In the absence of direct measurements, the distribution must be tomographically reconstructed from its projections. In this paper, we highlight that such problems can be severely underdetermined and that entropy maximization is the most conservative solution strategy. We leverage normalizing flows—invertible generative models—to extend maximum-entropy tomography to 6D phase space and perform numerical experiments to validate the model performance. Our numerical experiments demonstrate that flow-based entropy estimates are consistent with 2D maximum-entropy solutions and that normalizing flows can fit complex 6D phase space distributions to large measurement sets in reasonable time.
@@ I Introduction
Accelerators generate charged particle beams with tailored distributions in position-momentum space (phase space). Measuring the phase space distribution enables model-based beam optimization and control, but such measurements are challenging when the phase space is high-dimensional. If direct measurements [1, 2, 3] are unavailable, the distribution must be tomographically reconstructed from its projections. Fig. 1 shows a typical setup, where each projection corresponds to the beam’s spatial distribution after transport through an accelerator lattice (a series of electromagnetic focusing elements, such as dipole or quadrupole magnets). In the most challenging versions of the problem, we face a 4D or 6D reconstruction from 1D or 2D projections and (potentially) nonlinear transformations [4, 5, 6, 7, 8, 9, 10, 11, 12].
Figure 1: Generic phase space tomography setup. The $n$-dimensional phase space distribution $\rho(\bm{\mathbf{x}})$ travels through an accelerator lattice—a series of electromagnetic focusing elements—represented by a symplectic mapping $\bm{\mathbf{u}}_{k}=\mathcal{M}_{k}(\bm{\mathbf{x}})$. We measure the $m$-dimensional projections $g(\bm{\mathbf{u}}_{k_{\parallel}})$ in the transformed space and use this data to reconstruct $\rho(\mathbf{x})$. If the lattice contains nonlinear elements, the projections correspond to curved integration surfaces in the original phase space.
@@
It is possible for 6D reconstructions to be severely ill-posed. Since the measured dimension is fixed, the set of feasible distributions—those consistent with the measurements—may proliferate with the phase space dimension. The total number of views can be limited to tens, compared to hundreds in industrial or medical applications. We generally do not know the information-maximizing set of 6D phase space transformations under given measurement conditions, and accelerator constraints place many transformations out of reach. And we do not know the level to which modeling or measurement errors propagate through the 6D reconstruction.
@@
Entropy maximization is a conservative approach to ill-posed inverse problems and is thus well suited to high-dimensional tomography. [^1] Entropy maximization can be derived from self-consistency requirements on any strategy to infer probability distributions from incomplete data [13, 14]. The maximum-entropy distribution has minimal informational content, eliminating spurious correlations and deviating from prior knowledge only to the extent required by the data [15]. In this paper, we leverage normalizing flows—invertible generative models—to extend maximum-entropy tomography to 6D phase space. Our approach is a straightforward merger of two previous studies. Loaiza-Ganem, Gao, and Cunningham [16], first proposed the use of flows for entropy maximization subject to statistical moment constraints; we incorporate projection constraints using the differentiable physics simulations and projected density estimation proposed by Roussel et al. [11] in the Generative Phase Space Reconstruction (GPSR) framework. We refer to the resulting approach as MENT-Flow.
@@
We begin by deriving the form of the $n$-dimensional maximum-entropy distribution subject to $m$-dimensional projection constraints, extending the 2D and 4D analysis in [17, 7]. We then discuss the shortcomings of existing maximum-entropy tomography algorithms when $n=6$ and describe the flow-based solution. Finally, we perform numerical experiments to validate the model’s reliability in 2D settings and examine the effects of entropic regularization in 6D tomography.
@@ II Maximum entropy tomography
Let $\rho(\bm{\mathbf{x}})$ be a probability distribution over phase space coordinates $\bm{\mathbf{x}}\in\mathbb{R}^{n}$. We assume the $k$th measurement occurs after a symplectic transformation $\mathcal{M}_{k}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Splitting the transformed coordinates
@@
into a projection axis $\bm{\mathbf{u}}_{k_{\parallel}}\in\mathbb{R}^{m}$ and orthogonal integration axis $\bm{\mathbf{u}}_{k_{\perp}}\in\mathbb{R}^{n-m}$, we write the constraints as
@@
for measured projection $g_{k}(\bm{\mathbf{u}}_{k_{\parallel}})$. Letting prior $\rho_{*}(\bm{\mathbf{x}})$ encode our knowledge of the distribution in the absence of data, we aim to maximize the entropy
@@
subject to Eq. (2). We will refer to this problem as an $n$:$m$ reconstruction.
We can derive the form of the maximum-entropy distribution from a new functional
[…]
where $\lambda_{k}(\bm{\mathbf{u}}_{k_{\parallel}})$ are unknown functions [18]. Enforcing zero variation of $\Psi$ with respect to $\rho(\bm{\mathbf{x}})$ and $\lambda_{k}(\bm{\mathbf{u}}_{k_{\parallel}})$ gives
@@
The constrained maximum-entropy distribution function at point $\bm{\mathbf{x}}$ is the product of two terms: the prior distribution function at $\bm{\mathbf{x}}$ and the product of the positive Lagrange functions $h_{k}$ evaluated at the corresponding points $\bm{\mathbf{u}}_{k_{\parallel}}(\bm{\mathbf{x}})$ on the measurement axes. [^2] In MENT literature, the Lagrange multipliers/functions often refer to $\lambda_{k}$ in Eq. (5), but here we refer to the exponential of $\lambda_{k}$. Substituting Eq. (5) into Eq. (2) generates a set of coupled nonlinear integral equations from which $h_{k}$ are to be solved.
@@ II.1 MENT
$h_{k}(\bm{\mathbf{u}}_{k_{\parallel}})=\begin{cases}1,&\text{if}\ g_{k}(\bm{\mathbf{u}}_{k_{\parallel}})>0\\ 0,&\text{otherwise}\end{cases}.$ (6)
@@
$\tilde{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}})$ $=\int{\rho_{*}\left(\mathcal{M}_{k}^{-1}(\bm{\mathbf{u}}_{k})\right)\prod\limits_{j}h_{j}(\bm{\mathbf{u}}_{j_{\parallel}})d\bm{\mathbf{u}}_{k_{\perp}}}$ (8)
$=h_{k}(\bm{\mathbf{u}}_{k_{\parallel}})\int{\rho_{*}\left(\mathcal{M}_{k}^{-1}(\bm{\mathbf{u}}_{k})\right)\prod\limits_{j\neq k}h_{j}(\bm{\mathbf{u}}_{j_{\parallel}})d\bm{\mathbf{u}}_{k_{\perp}}}$
@@
MENT is a powerful tomographic algorithm. Its unique advantage is that it maximizes entropy by design: fitting the data generates an exact solution to the constrained optimization problem. MENT is also efficient (it stores the exact number of parameters needed to define the maximum-entropy distribution and typically converges in a few epochs) and essentially hyperparameter-free. Additionally, although not yet reported in the literature, MENT does not assume the phase space transformations are linear.
We showed above that MENT extends to $n$:$m$ tomography. In practice, the value of $n$ is limited by the integrals required to compute the distribution’s projections. Ongoing work aims to demonstrate efficient implementations when $n=4$ [7, 20]. Extension to $n=6$ may be possible, but it has not been demonstrated, and the runtime may be slow if there are many high-resolution measurements. Additionally, even if the algorithm converged, generating particles would require sampling from an expensive-to-evaluate 6D distribution function.
@@ II.2 MENT-Flow
$L[\rho(\bm{\mathbf{x}}),\rho_{*}(\bm{\mathbf{x}}),\{{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}})\}]=-H[\rho(\bm{\mathbf{x}}),\rho_{*}(\bm{\mathbf{x}})]+\mu\sum_{k}{D[{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}}),\tilde{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}})]}$ (9)
@@
for an increasing sequence of penalty parameters $\mu$. Here, $D[{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}}),\tilde{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}})]$ is a non-negative number quantifying the discrepancy between the measured and simulated projections, which we choose to be the Kullback-Leibler (KL) divergence. Although exact solutions may require $\mu\rightarrow\infty$, approximate solutions obtained with finite $\mu$ are often sufficient. [^3] More sophisticated Augmented Lagrangian (AL) methods [16, 21] are available to stabilize the optimization and avoid large penalty parameters, but we did not see significant improvements using AL in our numerical experiments. This approach requires a parameterization of $\rho(\bm{\mathbf{x}})$. Direct parameterizations require 6D integration or Monte Carlo sampling to compute the distribution’s entropy and projections. Indirect parameterizations instead define
$\bm{\mathbf{x}}=\mathcal{F}(\bm{\mathbf{z}}),$ (10)
@@
where $\mathbf{z}$ is drawn from a base distribution $\rho_{0}(\bm{\mathbf{z}})$ defined in a normalized space. Typically, $\rho_{0}(\bm{\mathbf{z}})$ is an easy-to-sample-from distribution such as a standard Gaussian distribution. Roussel et al. [11] adopted this approach, using a neural network to represent $\mathcal{F}$ and using differentiable beam dynamics simulations and kernel density estimation to fit the network parameters to projection data. However, it is difficult to maximize the entropy without access to $\rho(\bm{\mathbf{x}})$. [^4] Roussel et al. [11] proposed to maximize the emittance, or root-mean-square (rms) volume, $\varepsilon={|\bm{\mathbf{\Sigma}}|}^{1/2}$, where $\bm{\mathbf{\Sigma}}=\langle\bm{\mathbf{x}}\bm{\mathbf{x}}^{T}\rangle$ is the $n\times n$ covariance matrix of second order moments, as a proxy for the entropy. For certain distributions, the logarithm of the emittance is proportional to the entropy, but this is not true in general. Like entropy maximization, emittance maximization removes unnecessary linear correlations from the reconstructed distribution. However, it cannot remove nonlinear correlations, as the emittance depends only on second-order moments. Furthermore, the maximum-emittance distribution is not unique. In linear systems, the covariance matrix is typically overdetermined by the tomographic measurements, i.e., all distributions that fit the data have the same emittance. Particle-based entropy estimates based on k nearest neighbors [22] may perform better.
Normalizing flows are indirect parameterizations that provide access to the probability density. Flows define $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ as an invertible, differentiable map and compute the density from
@@
is the $n\times n$ Jacobian matrix of $\mathcal{F}$. Given $N$ flow samples $\{\bm{\mathbf{x}}_{i}\}$, an unbiased entropy estimate is available: [16]:
$-H[\rho(\bm{\mathbf{x}}),\rho_{*}(\bm{\mathbf{x}})]$ $=\mathbb{E}_{\rho(\bm{\mathbf{x}})}\left[\log\rho(\bm{\mathbf{x}})-\log\rho_{*}(\bm{\mathbf{x}})\right]$ (13)
$\approx\frac{1}{N}\sum_{i=1}^{N}\left(\log\rho(\bm{\mathbf{x}}_{i})-\log\rho_{*}(\bm{\mathbf{x}}_{i})\right)$
Since $\mathcal{F}$ is differentiable, we can maximize the entropy estimate via, e.g., stochastic gradient descent [16]. We call this approach MENT-Flow when combined with differentiable tomographic constraints.
@@
It is not immediately obvious whether flows can learn complex 6D phase space distributions from projection data in reasonable time. Flows are restricted to invertible transformations that preserve the topological features of the base distribution; for example, flows cannot perfectly represent disconnected modes if the base distribution has a single mode [23]. Building complex flows requires layering transformations, either as a series of maps (discrete flows) or a system of differential equations (continuous flows), often leading to large models and expensive training. [^5] A relevant example comes from Green, Ting, and Kamdar [24], who used continuous flows for 6D phase space density estimation from measured stellar phase space coordinates; training times ranged from hours to days on a GPU, depending on the distribution complexity, with approximately $10^{4}$ particles per batch.
@@
We found that Neural Spline Flows (NSF) [25] provide a sufficient blend of speed and power. Each flow layer is an autoregressive transformation of the form
$z_{i}\rightarrow\mathcal{F}(z_{i};c_{i}(z_{1},\dots,z_{i-1})),$ (14)
where $\bm{\mathbf{z}}=[z_{1},\dots,z_{n}]^{T}$. Here, $\mathcal{F}$ is an invertible function conditioned on $c_{i}(z_{1},\dots,z_{i-1})$. The conditioner generates a triangular Jacobian whose determinant can be efficiently computed. We implemented the conditioner using a masked neural network [26] and the transformer using rational-quadratic splines [25]. Representational power increases with the number of neural network parameters, the number of knots in the spline transformer, and the number of flow layers. For the composition of $T$ layers
@@
$\left|J_{\mathcal{F}}(\bm{\mathbf{z}}_{0})\right|=\prod_{t=1}^{T}{\left|J_{\mathcal{F}_{t}}(\bm{\mathbf{z}}_{t-1})\right|.}$ (17)
MENT-Flow increases the reconstruction model complexity and does not guarantee an exact entropy maximum. However, MENT-Flow scales straightforwardly to $n$-dimensional phase space and immediately generates independent and identically distributed samples from the reconstructed distribution.

### v3

@@ Abstract
Particle accelerators generate charged particle beams with tailored distributions in six-dimensional position-momentum space (phase space). Knowledge of the phase space distribution enables model-based beam optimization and control. In the absence of direct measurements, the distribution must be tomographically reconstructed from its projections. In this paper, we highlight that such problems can be severely underdetermined and that entropy maximization is the most conservative solution strategy. We leverage normalizing flows—invertible generative models—to extend maximum-entropy tomography to six-dimensional phase space and perform numerical experiments to validate the model’s performance. Our numerical experiments demonstrate consistency with exact two-dimensional maximum-entropy solutions and the ability to fit complicated six-dimensional distributions to large measurement sets in reasonable time.
@@ I Introduction
Particle accelerators generate charged particle beams with tailored distributions in position-momentum space (phase space). Measuring the phase space distribution in the accelerator enables model-based beam optimization and control and provides a valuable benchmark for simulation codes. In the absence of direct measurements [1, 2, 3], the distribution must be reconstructed from its projections. [^1] The 1D beam density can be measured by recording the secondary electron emission from a wire swept across the beam. Scintillating screens provide 2D projections of electron beams or low-intensity, low-energy hadron beams. 2D projections of higher energy hadron beams are only available from specialized diagnostics such as laser wires [4]. Fig. 1 illustrates a generic setup in which the beam is measured under varying accelerator conditions and reconstructed at a location upstream of the measurement device.
Figure 1: Generic phase space tomography setup. An initial phase space distribution $\rho(\bm{\mathbf{x}})$ travels through an accelerator segment represented by the symplectic transformation $\bm{\mathbf{u}}_{k}=\mathcal{M}_{k}(\bm{\mathbf{x}})$ for measurement index $k$. Each projection $g_{k}(\bm{\mathbf{u}}_{k_{\parallel}})$ of the transformed distribution is a different low-dimensional view of the initial distribution.
If the accelerator linearly transforms the phase space coordinates and does not couple the three planes of motion, one can reconstruct the 2D phase space distribution using conventional tomography algorithms. It is more challenging to reconstruct the 4D or 6D phase space distribution. Many conventional algorithms represent the distribution on a grid and face massive storage requirements as the phase space dimension scales [5]. Several authors have developed new algorithms and diagnostics to sidestep this issue and fit 4D phase space distributions to 2D projections [6, 7, 5, 8, 9, 10]. There has also been one extension to 5D phase space [11], and one extension to 6D phase space [12].
@@
An additional challenge is that high-dimensional reconstructions may be ill-posed; since the measured dimension is fixed, the set of feasible distributions (those consistent with the measurements) may proliferate with the phase space dimension. It is usually infeasible to compensate by exponentially increasing the number of measurements, as one is typically limited to tens of views because of slow diagnostic devices and limited beam time. Additionally, it is not yet clear how to derive the information-maximizing set of high-dimensional phase space transformations under given measurement conditions—and in any case, accelerator constraints place many transformations out of reach.
To select a single solution from the feasible set, our strategy is to define a prior probability distribution over the phase space coordinates and update the prior to a posterior by incorporating the information in the measurements. Our information comes in the form of constraints, and we perform the update by maximizing a convex functional subject to these constraints. Under basic self-consistency requirements, the functional must be the relative entropy [13, 14, 15, 16]. Entropy maximization ensures that the posterior does not deviate from the prior unless forced to by the data. This is a conservative strategy that eliminates all spurious features from the reconstructed distribution.
@@
Entropy maximization is not always feasible, especially in high dimensions, because it entails a highly nonlinear constrained optimization. Although a reliable exact maximum-entropy algorithm exists for 2D tomography, its computational complexity scales exponentially with the phase space dimension, rendering its extension to 6D prohibitively expensive at this time. In this paper, we leverage normalizing flows—invertible generative models—to find approximate 6D maximum-entropy solutions. Our approach is a straightforward extension of two previous studies. Loaiza-Ganem, Gao, and Cunningham [17] first proposed the use of normalizing flows for entropy maximization subject to statistical moment constraints; we incorporate projection constraints using the differentiable physics simulations and projected density estimation proposed by Roussel et al. [10] in the Generative Phase Space Reconstruction (GPSR) framework. We refer to the resulting approach as MENT-Flow.
@@
We begin by deriving the form of the $n$-dimensional maximum-entropy distribution subject to $m$-dimensional projection constraints, following the analysis in [18]. We then discuss the shortcomings of existing maximum-entropy tomography algorithms when $n=6$ and describe the flow-based solution. Finally, we perform numerical experiments to validate the model’s reliability in 2D settings and examine the effects of entropic regularization in 6D tomography.
@@ II Maximum entropy tomography
Let $\rho_{*}(\bm{\mathbf{x}})$ be a prior probability distribution over the phase space coordinates $\bm{\mathbf{x}}\in\mathbb{R}^{n}$. We wish to update the prior to a posterior $\rho(\bm{\mathbf{x}})$ by maximizing the relative entropy
[…]
while enforcing consistency with a set of $m$-dimensional projections. We will refer to this problem as an $n$:$m$ reconstruction.
We assume the $k$th measurement occurs after a symplectic transformation $\mathcal{M}_{k}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. By splitting the transformed coordinates
@@
into a projection axis $\bm{\mathbf{u}}_{k_{\parallel}}\in\mathbb{R}^{m}$ and orthogonal integration axis $\bm{\mathbf{u}}_{k_{\perp}}\in\mathbb{R}^{n-m}$, we can write the constraints as
$G_{k}\left[\rho(\bm{\mathbf{x}})\right]=g_{k}(\bm{\mathbf{u}}_{k_{\parallel}})-\tilde{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}})=0,$ (3)
where $g_{k}(\bm{\mathbf{u}}_{k_{\parallel}})$ are the measured projections and
@@
are the simulated projections. The form of the maximum-entropy posterior distribution can be derived from a new functional
@@
where $\lambda_{k}(\bm{\mathbf{u}}_{k_{\parallel}})$ are Lagrange multipliers [19]. Enforcing zero variation of $\Psi$ with respect to $\rho(\bm{\mathbf{x}})$ and $\lambda_{k}(\bm{\mathbf{u}}_{k_{\parallel}})$ gives
@@
where we have defined $h_{k}(\bm{\mathbf{u}}_{k})=\exp(\lambda_{k}(\bm{\mathbf{u}}_{k}))$. Substituting Eq. (6) into Eq. (3) generates a set of coupled nonlinear integral equations from which $h_{k}$ are to be solved.
@@ II.1 MENT
$h_{k}(\bm{\mathbf{u}}_{k_{\parallel}})=\begin{cases}1,&\text{if}\ g_{k}(\bm{\mathbf{u}}_{k_{\parallel}})>0\\ 0,&\text{otherwise}\end{cases},$ (7)
@@
$\tilde{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}})=\int{\rho_{*}(\bm{\mathbf{x}}(\bm{\mathbf{u}}_{k}))\prod\limits_{j}h_{j}(\bm{\mathbf{u}}_{j_{\parallel}}(\bm{\mathbf{u}}_{k}))d\bm{\mathbf{u}}_{k_{\perp}}}$ (9)
@@
MENT maximizes entropy by design: fitting the data generates an exact solution to the constrained optimization problem. MENT is also efficient: it stores the exact number of parameters needed to define the maximum-entropy distribution and typically converges in a few epochs. Finally, MENT is essentially free of hyperparameters.
The MENT formulation above is valid for $n$:$m$ tomography, but the integrals in Eq. (9) limit the value of $n$ in practice. Ongoing work aims to demonstrate efficient implementations when $n=4$ [18, 22]. Extension to $n=6$ may be possible, but it has yet to be demonstrated, and the runtime would likely be quite long if there were many high-resolution measurements. Even if the algorithm converged, sampling particles from the posterior (Eq. (6)) would be a nontrivial extra step.
@@ II.2 MENT-Flow
$L=-H[\rho(\bm{\mathbf{x}}),\rho_{*}(\bm{\mathbf{x}})]+\mu\sum_{k}{D[{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}}),\tilde{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}})]}$ (10)
@@
for an increasing sequence of penalty parameters $\mu$. Here, $D[{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}}),\tilde{g}_{k}(\bm{\mathbf{u}}_{k_{\parallel}})]$ is a non-negative number quantifying the discrepancy between the measured and simulated projections, which we choose to be the Kullback-Leibler (KL) divergence. Exact solutions may require $\mu\rightarrow\infty$, but approximate solutions obtained with finite $\mu$ are often sufficient.
The above approach requires us to represent the distribution using a finite set of parameters, $\bm{\mathbf{\theta}}$. Grid-based representations become expensive when $n\geq 4$. An attractive alternative is to directly predict the value of $\rho(\bm{\mathbf{x}})$ up to a normalization constant; however, computing the distribution’s entropy (Eq. (1)) and projections (Eq. (4)) would require expensive integration or Monte Carlo sampling. We might instead define the distribution indirectly via the transformation
$\bm{\mathbf{x}}=\mathcal{F}(\bm{\mathbf{z}};\bm{\mathbf{\theta}}),$ (11)
where $\mathcal{F}:\mathbb{R}^{n^{\prime}}\rightarrow\mathbb{R}^{n}$ is a map parameterized by $\bm{\mathbf{\theta}}$, and $\mathbf{z}\in\mathbb{R}^{n^{\prime}}$ is a random variable drawn from a base distribution $\rho_{0}(\bm{\mathbf{z}})$ defined in a “normalized” or “latent” space. The base distribution is typically a Gaussian. Sampling from $\rho(\bm{\mathbf{x}})$ reduces to sampling from $\rho_{0}(\bm{\mathbf{z}})$ and applying the unnormalizing transformation in Eq. (11). Thus, Eq. (11) defines a generative model. In most generative models, $\mathcal{F}$ is a neural network trained to learn an unknown distribution from data samples [23].
Roussel et al. [10] showed that generative models can also be trained to match projections of the unknown distribution. To train the model via gradient descent, the transformations from the base distribution to the measurement locations must be differentiable:
$\bm{\mathbf{u}}_{k}=\mathcal{M}_{k}(\mathcal{F}(\bm{\mathbf{z}};\bm{\mathbf{\theta}})).$ (12)
@@
This is possible using a differentiable beam physics simulation [24] to represent $\mathcal{M}_{k}$. The calculation of the projected density $\tilde{g}_{k}(\bm{\mathbf{u}}_{k})$ in Eq. (4) must also be differentiable. This is possible using 1D or 2D kernel density estimation. It is, however, difficult to maximize the entropy without access to the density $\rho(\bm{\mathbf{x}})$. [^2] Roussel et al. [10] proposed to maximize the emittance, or root-mean-square (rms) volume, $\varepsilon={|\bm{\mathbf{\Sigma}}|}^{1/2}$, where $\bm{\mathbf{\Sigma}}=\langle\bm{\mathbf{x}}\bm{\mathbf{x}}^{T}\rangle$ is the $n\times n$ covariance matrix of second order moments, as a proxy for the entropy. For certain distributions, the logarithm of the emittance is proportional to the entropy, but this is not true in general. Like entropy maximization, emittance maximization removes unnecessary linear correlations from the reconstructed distribution. However, it cannot remove nonlinear correlations, as the emittance depends only on second-order moments. Furthermore, the maximum-emittance distribution is not unique. In linear systems, the covariance matrix is typically overdetermined by the tomographic measurements, i.e., all distributions that fit the data have the same emittance. Particle-based entropy estimates based on k nearest neighbors [25] may perform better.
A normalizing flow, or simply flow, follows the same paradigm but provides access to the probability density. A normalizing flow is a differentiable map $\mathcal{F}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ with a differentiable inverse $\mathcal{F}^{-1}$ [26]. These properties ensure we can compute the change in probability density under the transformation in Eq. (11):
@@
is the $n\times n$ Jacobian matrix of $\mathcal{F}$, accounting for volume change, and $\bm{\mathbf{z}}=[z_{1},\dots,z_{n}]^{T}$. To compute the probability density at $\bm{\mathbf{x}}$, we flow backward and multiply the base distribution at $\bm{\mathbf{z}}$ by the absolute value of the Jacobian matrix determinant. To generate samples $\{\bm{\mathbf{x}}_{i}\}$, we sample points $\{\bm{\mathbf{z}}_{i}\}$ from the base distribution and unnormalize them by flowing forward. We also obtain the probability density $\{\rho(\bm{\mathbf{x}}_{i})\}$ at each sampled point by tracking the Jacobian matrix determinant during this forward pass.
The ability to generate particles and evaluate the probability density at each particle is useful for computing expected values. Given $N$ samples $\{\bm{\mathbf{x}}_{i}\}$ from $\rho(\bm{\mathbf{x}})$, the following expression is an unbiased estimate of the expected value of a functional $Q[\rho(\bm{\mathbf{x}})]$:
[…]
Since the entropy is the expected value of $\log(\rho(\bm{\mathbf{x}})/\rho_{*}(\bm{\mathbf{x}}))$, the following expression is an unbiased estimate of the entropy [17]:
$H[\rho(\bm{\mathbf{x}}),\rho_{*}(\bm{\mathbf{x}})]$ $\approx-\frac{1}{N}\sum_{i=1}^{N}\log(\rho(\bm{\mathbf{x}}_{i})/\rho_{*}(\bm{\mathbf{x}}_{i}))$ (16)
Since the estimate is differentiable, it can be maximized via stochastic gradient descent [17]. Thus, our approach is to use the Generative Phase Space Reconstruction (GPSR) method [10] with a normalizing flow instead of a conventional neural network. We call this approach MENT-Flow, in reference to the MENT algorithm.
@@
It is not immediately obvious whether normalizing flows can learn complex 6D distributions from projections in reasonable time. Flows preserve the topological features of the base distribution; for example, flows cannot perfectly represent disconnected modes if the base distribution has a single mode [27]. Thus, building complex flows requires layering transformations, either as a series of maps (discrete flows) or a system of differential equations (continuous flows), often leading to large models and expensive training. [^3] A relevant example comes from Green, Ting, and Kamdar [28], who used continuous flows for 6D phase space density estimation from measured stellar phase space coordinates. Training times ranged from hours to days on a GPU, depending on the distribution complexity, with approximately $10^{4}$ particles per batch.
@@
We found that neural spline flows (NSF) [29] provide a sufficient blend of speed and power. In this model, the unnormalizing transformation $\mathcal{F}$ has the following autoregressive form:
$x_{i}=\tau(z_{i};c_{i}(z_{1},\dots,z_{i-1})),$ (17)
where $\bm{\mathbf{x}}=[x_{1},\dots,x_{n}]^{T}$, $\bm{\mathbf{z}}=[z_{1},\dots,z_{n}]^{T}$, and $\tau$ is an invertible function parameterized by $c_{i}(z_{1},\dots,z_{i-1})$. The transformation is invertible for any $c_{i}$, and since $c_{i}$ depends only on the first $i-1$ dimensions, the transformation has a triangular Jacobian matrix whose determinant can be computed efficiently. In the NSF model, the 1D transformer ($\tau$) is a monotonic rational-quadratic spline [29]. The spline is defined by the locations of $K$ different knots and the derivative at each knot. These parameters are provided by the conditioner ($c$), a masked neural network [30] in which connections between nodes in a regular feedforward neural network are removed to produce the triangular Jacobian matrix.
The model’s representational power increases with the number of parameters in the masked neural network and the number of knots in the rational-quadratic splines. We can also define more than one flow layer. For the composition of $T$ layers
@@
$\left|\det J_{\mathcal{F}}(\bm{\mathbf{z}}_{0})\right|=\prod_{t=1}^{T}{\left|\det J_{\mathcal{F}_{t}}(\bm{\mathbf{z}}_{t-1})\right|.}$ (20)
Compared to MENT, MENT-Flow increases the reconstruction model complexity and does not guarantee an exact entropy maximum. However, MENT-Flow scales straightforwardly to $n$-dimensional phase space and immediately generates independent and identically distributed samples from the reconstructed distribution function.